OpenStudy (swaf11):

8 1/3 *s s=4 1/2 a.)37 1/2 b.)32 1/5 c.)1 1/6

jabez177 (jabez177):

First step, we need to convert the Whole fractions to Improper fractions. 8 x 3 + 1 = ?

OpenStudy (swaf11):

25/3

OpenStudy (swaf11):

is 8 1/3 improper, right?

jabez177 (jabez177):

Now 4 x 2 + 1 = ?

OpenStudy (swaf11):

9/2

jabez177 (jabez177):

No, \[8\frac{ 1 }{ 3 }\] is the whole fraction.

OpenStudy (swaf11):

ohhh ok

jabez177 (jabez177):

And nice, it's 9/1 :)

jabez177 (jabez177):

I mean 9/2

OpenStudy (swaf11):

so u multiply 8 1/3 by 9/2

jabez177 (jabez177):

Now, \[\frac{ 25 }{ 3 }\times \frac{ 9 }{ 2 }\]

jabez177 (jabez177):

How many times can 3 go into 9? Meaning, 3 x what = 9?

OpenStudy (swaf11):

3

OpenStudy (swaf11):

wait how does 3*3 relate to 25/3*9/2?

jabez177 (jabez177):

Good so the 3 turns into 1 and the 9 turns into 3. So now it is \[\frac{ 25 }{ 1 }\times \frac{ 3 }{ 2 }\]

OpenStudy (swaf11):

oh ok

jabez177 (jabez177):

So now what is 25 x 3?

OpenStudy (swaf11):

75

jabez177 (jabez177):

1 x 2 = ?

OpenStudy (swaf11):

2

jabez177 (jabez177):

Good. Now 75 divided by 2 = what?

OpenStudy (swaf11):

37.5

jabez177 (jabez177):